OpenStudy (anonymous):

Rework problem 1 in section 4.1 of your text, involving the flipping of a loaded coin, but assume that Pr[H] = 0.3. Also, assume that the coin is flipped 4 times, and the random variable X is defined to be 3 times the number of heads minus 2 times the number of tails. How many different values are possible for the random variable X?

OpenStudy (anonymous):

x 3*0-2*4=-8 3*1-2*3=-3 3*2-2*2=2 3*3-2*1=7 3*4-2*0=12

OpenStudy (anonymous):

So, how many different values are possible for the random variable X?